OpenStudy (anonymous):

ok i need help with this word problem "The sum of the digits of a two-digit number is 8. If 16 is added to the original number, the result is 3 times the original number with its digits reversed. Find the original number." im pretty sure i have it but i dont know what variable to use for the 2 digit number

OpenStudy (anonymous):

Le the digits be x and y. Hence the original number is = 10x + y. Since the digits add up to 8. Hence, x + y = 8 Again adding 16 to the original number we get - 10x + y + 16 This number is 3 times of the original number reversed 3(10y + x) Hence, 10x + y + 16 = 3 (10y +x) => 7x - 29y = -16 Replacing y = 8 - x, we have 7x - 29(8 - x) = -16 => x = 6, hence y = 8 - 6 = 2 Hence the original number is 10 * 6 + 2 = 62

OpenStudy (anonymous):

where did the -10 come from?

OpenStudy (anonymous):

no no :) that is not -10, I just written "we get -' the number actually is 10x + y + 16

OpenStudy (anonymous):

haha sorry then where did the 10 come from

OpenStudy (mertsj):

x is the ten's digit. It's value is 10x

OpenStudy (anonymous):

ohh ok thanks guys

OpenStudy (anonymous):

im gonna have more of these coming so get ready haha:)